A ray of light of wavelength 36.5x10^-8 m and intensity of 10^-8 Wm^-2 is incident perpendicularly on a surface. Absorption coefficient of that surface is 0.8 and work function is 1.6eV. Calculate

(a) rate of production of electron from unit area

(b) absorbed energy per square meter

(c) kinetic energy of the produced electrons


1
Expert's answer
2021-09-13T11:00:56-0400

Given:

"\lambda=36.5*10^{-8}\:\rm m"

"I=10^{-8}\:\rm W/m^2"

"k=0.8"

"\phi=1.6\:\rm eV=2.56*10^{-19}\:\rm J"


(a) the rate of production of electron from unit area is given by

"\frac{dN}{Adt}=\frac{kI}{E_0}=\frac{kI\lambda}{hc}"

"\frac{dN}{Adt}=\frac{0.8*10^{-8}*36.5*10^{-8}}{6.63*10^{-34}*3.0*10^{8}}=1.47*10^{10} \:\rm 1/m^2\cdot s"

(b) the absorbed energy per square meter is given by

"E/At=kI=0.8*10^{-8}=8.0*10^{-9}\:\rm W/m^2"

(c) the kinetic energy of the produced electrons is given by

"E_k=hc/\lambda-\phi"

"E_k=\frac{6.63*10^{-34}*3.0*10^{8}}{36.5*10^{-8}}-2.56*10^{-19}\\
=2.88*10^{-19} \:\rm J=1.8\:\rm eV"